# Fields of The World: A Machine Learning Benchmark Dataset For Global Agricultural Field Boundary Segmentation

### v1

@@ 1 Introduction
We include training, validation, and test sets for each country, drawing from existing splits where possible to maximize compatibility with existing work. We propose benchmark tasks that mimic real-world scenarios relevant to downstream users of ML-digitized field boundary datasets (e.g., region-specific evaluation, transfer learning, and zero-shot generalization). Finally, we perform experiments to demonstrate the value of the FTW dataset and provide baseline results for benchmark tasks. We release all code via Github, data via Source Cooperative, and data loaders and pre-trained models via TorchGeo for Fields of The World.
@@ 2 Dataset Description
We tiled each sample grid into $1536\times 1536$ meter sample patch ROIs (regions of interest), which we refer to as ‘chips’. For India and Rwanda, we created a $1536\times 1536$ meter chip ROI around the center of each label cluster.
@@
We converted the label polygon datasets to the fiboa specification (fiboa contributors 2024). If a dataset was sub-sampled from a larger dataset, we only converted the sub-sampled version. Per the fiboa core specification, each dataset has per-polygon attributes id (unique field identifier), determination_datetime (last timestamp at which the field boundary existed/was observed), area (field area in hectares), and geometry (field polygon geometry; we use the WGS84/EPSG:4326 coordinate reference system). If crop type or other attributes were available from the source dataset, we included those additional attributes. Converted GeoParquet files are available on Source Cooperative at https://beta.source.coop/repositories/kerner-lab/fields-of-the-world/. The README for each dataset provides details including the source dataset license (extends to the FTW subset using it) and source data link. We also provide GeoParquet files for the sample grids and sample chip ROIs.
@@
For each country, we collected images from two date ranges (Table 1). Growing seasons vary greatly around the world (and even within some countries). To choose the date ranges for each country, we looked at the country’s crop calendar which specifies the planting, mid-season, harvesting, and off-season months for the main crops (USDA Foreign Agricultural Service 2024). We inspected the available satellite images in the sample area(s) in two date ranges spanning the planting/mid-season and harvesting/off-season months. If a country had multiple growing seasons (e.g., winter and summer crops), we visualized both seasons and chose the one appearing most active. We then iteratively adjusted the date ranges to account for good contrast between images and cloud cover. After adjustment, the date ranges do not necessarily match the growing season stages, so we call them Window A and B.
@@
We resized each chip to $256\times 256$ pixels. Each chip is stored as a GeoTIFF with EPSG:4326 coordinate reference system. We normalize images during training by dividing each channel by $3000$ (an approximate mean value).
@@ 3 Dataset Analysis and Related Work
Figure 3: The distribution of (log) field area across four countries (top) and three datasets (bottom).
The dramatic differences in field morphology across the globe motivated the construction of FTW (Figure 1). FTW has significant advantages over previous datasets for field instance segmentation in terms of (i) geographic representation and extent, (ii) annotation volume, and (iii) annotation/scene complexity. The most relevant datasets for comparison are AI4Boundaries (d’Andrimont et al. 2023), AI4SmallFarms (Persello et al. 2023), PASTIS (Sainte Fare Garnot and Landrieu 2021), and PASTIS-R (Garnot, Landrieu, and Chehata 2022). We only include datasets that explicitly label individual field instances, excluding datasets that only include agriculture-related segmentation labels since field instances enable a broader range of applications. We also exclude datasets that provide field polygons but no imagery, such as the field boundary dataset created by the French Land Parcel Identification System (The Service and Payment Agency (2024) ASP). The lack of standardized satellite imagery for polygon-only datasets hampers the creation and comparison of automated algorithms for field boundary delineation and related tasks.
@@
The FTW dataset captures greater morphological diversity of agricultural field instances than any other dataset. Figure 1 shows example field boundaries from four countries. The diverse shapes and sizes reflect the unique combination of topographical, environmental, and historical factors that influenced the development of field boundaries in each region. Figure 3 (top) shows the dramatic difference between the distribution of field area of different countries. There is little overlap between the distributions of Vietnam (small field sizes) and Brazil (large field sizes). Figure 3 (bottom) shows the distribution field areas for AI4SmallFarms, AI4Boundaries, and FTW (note that we did not include PASTIS or PASTIS-R since both datasets provide field boundaries in raster, not vector, format so we could not compute morphological statistics). This highlights that AI4SmallFarms consists mostly of small-area fields, AI4Boundaries consists mostly of medium-area fields, and FTW has a broader distribution of field areas. There is also significant diversity in terms of field shape complexity (see visualizations in supplement). For example, Estonia and Slovakia have complex field shapes, with an average number of polygon vertices of 111.9 and 95.4, respectively. In contrast, Kenya and Rwanda have simpler field shapes, with an average number of vertices of 4.5 and 5.6, respectively.
@@ 4 Baseline Experiments
A common approach to field instance segmentation is binary segmentation followed by polygonization of predicted raster masks (Persello et al. 2023). We follow this approach in our experiments. Unless specified otherwise, we used a U-net with EfficientNet-b3 backbone with inputs consisting of concatenated 4-channel RGB-NIR images from both Window A and B (in that order). We initialized the RGB channels using ImageNet RGB weights and NIR channels using random weights. We optimized cross-entropy loss with class weights inversely proportional to each class’s frequency in the training set for each experiment. We trained all models for 100 epochs. We used a fixed random seed for all experiments (randomly chosen). We did not perform any further hyperparameter tuning. Experiments required 4 A100 and 8 V100 GPUs for approximately one week.
We evaluate models using semantic (pixel-level) and instance (object-level) segmentation metrics. We provide evaluation functions in the FTW code repository for pixel-level intersection over union (IoU), precision, and recall, and object-level precision and recall. We use rasterio to convert binary segmentation masks to polygons then computed object-level metrics using an IoU threshold of $0.5$.
@@
Fields of The World includes multiple target mask formats and satellite images from two dates and four spectral channels, giving users many modeling choices. In the experimental setup of semantic segmentation followed by polygonization, there are also many modeling choices such as which architecture to choose. We performed several experiments to assess the impact of these choices on model performance. In each experiment, we evaluate performance using the test set for each country. We report results for countries with presence/absence labels spanning a range of field sizes: Slovenia - SVN (average field size $0.64$ ha), France - FRA (average $5.7$ ha), and South Africa - ZAF (average $13.8$ ha). We provide results for all countries in the supplementary material. An extensive search of modeling configurations is beyond the scope of this work, but we hope future studies will explore a greater range of options using the FTW dataset.
@@
We evaluated the two types of target masks provided in FTW: binary field extent (field interior vs. background) and 3-class (field interior, boundary, and background) masks. In presence-only countries, pixels with an unknown class are labeled as background. We evaluated two scenarios: 1) when computing the loss, ignore all pixels labeled background for presence-only countries, and 2) compute the loss for all pixels treating unknown labels as background. Before computing test metrics, we converted outputs from all models to binary field extent masks to ensure a common evaluation basis.
Rows 1 and 3 of Table 3 show that almost all metrics are higher when training with 3-class masks than binary masks. Rows 2 and 4 compare when unknown-label pixels are counted as background or ignored for presence-only samples when training with 3-class masks. Although object precision is lower when ignoring unknown pixels, object recall is significantly higher, especially for Slovenia. Figure 4 shows that the segmentation masks predicted by both models are good, but 3-class masks improve the delineation of contiguous fields and thus object recall. From these results, we conclude that training with 3-class masks and ignoring presence-only background pixels is the strategy most likely to give good performance across all regions and use this setup for subsequent experiments.
@@
Some countries have many field boundary labels while others have few (Table 1). Previous work showed that performance on a data-scarce target region could be improved by pre-training models on a country with a large labeled dataset and then fine-tuning on the target region. Wang, Waldner, and Lobell (2022) fine-tuned a model for India after pre-training on France. Persello et al. (2023) fine-tuned for Cambodia and Vietnam after pre-training on the Netherlands.
@@
Motivated by the impressive performance of FTW pre-trained models without fine-tuning in Table 5, we used a FTW pre-trained model to predict field boundaries in Ethiopia (see Figure 5). This is a challenging region that is not part of FTW. The results show good qualitative performance that could be further improved with fine-tuning to local data and post-processing. This demonstrates the high potential of FTW to immediately be used in practice with little adaptation effort.
@@ 5 Discussion and Conclusion
ML research on automatic extraction of agricultural field boundaries from remotely sensed imagery is currently limited by a lack of ML-ready datasets and benchmarks to train and evaluate models on the global diversity of crop fields. These datasets are urgently needed in many applications for agriculture, climate change, and development. We designed Fields of The World to enable improved performance of ML models for field boundary segmentation in diverse global agricultural landscapes and enable granular country-scale evaluation for more countries than any existing dataset. We performed experiments to establish a performance baseline for the new FTW benchmark and showed that models trained with FTW improve performance over more geographically limited datasets analogous to existing benchmark datasets.
An exhaustive modeling search or proposal of novel field instance segmentation methods is beyond the scope of this work. Future work could build on the baselines we established by including more model architectures, including instance segmentation architectures (e.g., Mask-RCNN (He et al. 2017) or SAM (Kirillov et al. 2023)). We also did not evaluate self-supervised models (e.g., SatMAE (Cong et al. 2022) or Presto (Tseng et al. 2023)). Future work could also explore other methods of constructing target masks, motivated by our result that training with 3-class masks gave better performance than 2-class masks.
We provide complete metadata for the sample grids, sample chips, and field boundary polygons to enable future extensions of Fields of The World. For example, future work could add more spectral channels or sensors, timesteps, or sample locations. FTW can be extended as field polygons become available for more countries. We hope the community will build on FTW as research on this important task accelerates.
@@ Appendix A Supplementary Information
Figures 14 and 15 illustrate these predictions for a sample of 8 countries, representing both Presence/Absence and Presence-only regions (respectively). The heatmaps use three classes for prediction: Red for the Background class, Green for the Field Extent (Interior) class, and Blue for the Boundary class.
@@
The results in Tables 3, 4, and 5 of the main paper were run with one arbitrarily-chosen random seed. In supplement Tables 13-14, we report the average results across three random seeds for all test countries for the mask type experiment (Table 3 in the main paper). The standard deviation across random seeds for each experiment and test country are very small (0 or close to 0 for most metrics).

### v2

@@ 1 Introduction
We include training, validation, and test sets for each country, using existing splits when possible to maximize compatibility with existing work. We propose benchmark tasks that mimic real-world scenarios relevant to downstream users of field boundary datasets (e.g., region-specific evaluation, transfer learning, and zero-shot generalization). Finally, we perform experiments to demonstrate the value of the FTW dataset and provide baseline results for benchmark tasks. We release code via Github, data via Source Cooperative, and data loaders and pre-trained models via TorchGeo.
@@ 2 Dataset Description
We tiled each sample grid into $1536\text{m}\times 1536\text{m}$ sample patch ROIs (regions of interest), which we call ‘chips’. For India and Rwanda, we created a $1536\text{m}\times 1536\text{m}$ chip around the center of each label cluster.
@@
We converted the label polygon datasets to the fiboa specification (fiboa contributors 2024). If a dataset was sub-sampled from a larger dataset, we only converted the sub-sampled version. Per the fiboa core specification, each dataset has per-polygon attributes id (unique field identifier), determination_datetime (last timestamp at which the field boundary existed/was observed), area (field area in hectares), and geometry (field polygon geometry; we use the WGS84/EPSG:4326 coordinate reference system). We included crop type or other attributes when available. Converted GeoParquet files are available on Source Cooperative at https://beta.source.coop/repositories/kerner-lab/fields-of-the-world/. The README for each dataset provides details including the source dataset license (extends to the FTW subset using it) and link. We provide GeoParquet files for the sample grids and chip ROIs.
@@
For each country, we collected images from two date ranges (Table 1). Growing seasons vary greatly globally (and even within countries). To choose the date ranges, we looked at each country’s crop calendar which specifies the planting, mid-season, harvesting, and off-season months for the main crops (USDA Foreign Agricultural Service 2024). We inspected the available satellite images in the sample area(s) in two date ranges spanning the planting/mid-season and harvesting/off-season months. If a country had multiple growing seasons (e.g., winter and summer crops), we visualized both seasons and chose the one appearing most active. We then iteratively adjusted the date ranges to account for good contrast between images and cloud cover. After adjustment, the date ranges do not necessarily match the growing season stages, so we call them Window A and B.
@@
We resized each chip to $256\times 256$ pixels. Each chip is stored as a GeoTIFF with EPSG:4326. We normalize images during training by dividing each channel by $3000$ (an approximate mean value).
@@ 3 Dataset Analysis and Related Work
Figure 3: Field area distribution across four countries.
The dramatic differences in field morphology across the globe motivated the construction of FTW (Figure 1). FTW has significant advantages over previous field instance segmentation datasets in terms of (i) geographic representation and extent, (ii) annotation volume, and (iii) annotation/scene complexity. The most relevant datasets for comparison are AI4Boundaries (d’Andrimont et al. 2023), AI4SmallFarms (Persello et al. 2023), PASTIS (Sainte Fare Garnot and Landrieu 2021), and PASTIS-R (Garnot, Landrieu, and Chehata 2022). We only include datasets that explicitly label individual field instances, excluding semantic segmentation labels since field instances enable a broader range of applications. We also exclude datasets providing field polygons but no imagery, such as the field boundary dataset created by the French Land Parcel Identification System (The Service and Payment Agency (2024) ASP). The lack of standardized satellite imagery for polygon-only datasets hampers the creation and comparison of automated algorithms for field boundary delineation and related tasks.
@@
The FTW dataset captures greater morphological diversity of agricultural field instances than any other dataset. Figure 1 shows example field boundaries from four countries. The diverse shapes and sizes reflect the unique topographical, environmental, and historical factors that influenced the development of field boundaries in each region. Figure 3 shows the dramatic difference between the field areas in different countries. There is little overlap between the distributions of Vietnam (small fields) and Brazil (large fields). Figure 9 shows the distribution field areas for AI4SmallFarms, AI4Boundaries, and FTW (note that we did not include PASTIS/PASTIS-R since both provide field boundaries in raster, not vector, format so we could not compute morphological statistics). AI4SmallFarms consists mostly of small-area fields, AI4Boundaries consists mostly of medium-area fields, and FTW has a broader distribution of field areas. There is also significant diversity in field shape complexity (see visualizations in supplement). For example, Estonia and Slovakia have complex field shapes, with 111.9 and 95.4 polygon vertices on average, respectively. In contrast, Kenya and Rwanda have simpler field shapes, with 4.5 and 5.6 vertices on average, respectively.
@@ 4 Baseline Experiments
We follow the common approach to field instance segmentation of segmentation then polygonization of predicted raster masks (Persello et al. 2023). Unless specified otherwise, we used a U-net with EfficientNet-b3 backbone with inputs consisting of concatenated 4-channel RGB-NIR images from Window A and B. We found that this architecture performed well compared to other architectures and backbones (see Architectures paragraph in this section and Table 7 in supplement). We also found that concatenating both temporal windows and using four spectral bands performed best compared to other configurations (see Multi-temporal and multispectral channels paragraph and Table 4). We initialized the RGB channels using ImageNet and NIR channels using random weights. We optimized cross-entropy loss with class weights inversely proportional to each class’s frequency in the training set. We trained all models for 100 epochs. We used a fixed random seed for all experiments (randomly chosen). We did not perform hyperparameter tuning. Experiments required 4 A100 and 8 V100 GPUs for approximately one week.
We used semantic (pixel-level) and instance (object-level) segmentation metrics: pixel-level intersection over union (IoU), precision, and recall, and object-level precision and recall (functions in the FTW code repository). We converted segmentation masks to polygons using rasterio then computed object-level metrics with IoU threshold of $0.5$.
@@
Fields of The World includes multiple target mask formats and satellite images from two dates and four spectral channels, giving users many modeling choices. In semantic segmentation followed by polygonization, there are other choices such as which architecture to choose. We performed several experiments to assess the impact of these choices on model performance. In each experiment, we evaluate performance using the test set for each country. We report results for countries with presence/absence labels chosen to span a range of field sizes: Slovenia - SVN (average field size $0.64$ ha), France - FRA (average $5.7$ ha), and South Africa - ZAF (average $13.8$ ha). We provide results for all countries in the supplementary material. An extensive search of modeling configurations is beyond the scope of this work, but we hope future studies will explore a greater range of options using the FTW dataset.
@@
We evaluated the two types of target masks provided in FTW: 2-class (field interior vs. background) and 3-class (field interior, boundary, and background). In presence-only countries, pixels with an unknown class are labeled as background. We evaluated two scenarios: 1) when computing the loss, ignore all pixels labeled background for presence-only countries, and 2) compute the loss for all pixels treating unknown labels as background. Before computing test metrics, we converted outputs from all models to binary field extent masks to ensure a common evaluation basis.
Rows 1 and 3 of Table 3 show that almost all metrics are higher with 3-class masks than binary masks. Rows 2 and 4 compare when unknown-label pixels are counted as background or ignored for presence-only samples when training with 3-class masks. Although object precision is lower when ignoring unknown pixels, object recall is significantly higher, especially for Slovenia. Figure 4 shows that the segmentation masks predicted by both models are good, but 3-class masks improve the delineation of contiguous fields and thus object recall. From these results, we concluded that training with 3-class masks and ignoring presence-only background is most likely to give good performance across all regions and used this setup for subsequent experiments.
@@
Some countries have many labels while others have few (Table 1). Prior work showed that performance on a data-scarce region could be improved by pre-training models on a country with a large labeled dataset and then fine-tuning on the target region. Wang, Waldner, and Lobell (2022) fine-tuned a model for India after pre-training on France. Persello et al. (2023) fine-tuned for Cambodia and Vietnam after pre-training on the Netherlands.
@@
Motivated by the impressive performance of FTW pre-trained models without fine-tuning in Table 5, we used a FTW pre-trained model to predict field boundaries in Ethiopia, a challenging region not in FTW (Figure 5). The results show good qualitative performance that could be improved with local fine-tuning and post-processing. This shows the high potential of FTW to be immediately used in practice with little adaptation effort.
@@ 5 Discussion and Conclusion
ML research on automatic extraction of agricultural field boundaries from remotely sensed imagery is limited by a lack of ML-ready datasets to train and evaluate models on the global diversity of crop fields. These datasets are urgently needed in many applications for agriculture, climate change, and development. We designed Fields of The World to improve ML model performance for field boundary segmentation in diverse global agricultural landscapes and enable granular country-scale evaluation for more countries than any prior dataset. Our experiments established a performance baseline for the new FTW benchmark and showed that FTW-trained models perform better than more geographically limited datasets analogous to existing benchmarks.
Future work could build on these baselines by testing more model architectures, including instance segmentation architectures (e.g., Mask-RCNN (He et al. 2017) or SAM (Kirillov et al. 2023)) and geospatial foundation models (e.g., SatMAE (Cong et al. 2022) or Presto (Tseng et al. 2023)). Future work could also explore other methods of constructing target masks, motivated by our result that training with 3-class masks performed better than 2-class masks.
We provide complete metadata for sample grids, sample chips, and field boundary polygons to enable future extensions of Fields of The World. For example, future work could add more spectral channels or sensors, timesteps, or sample locations. FTW can be extended as field polygons become available for more countries. We hope the community will build on FTW as research on this important task grows.
Benchmarking on FTW
We hope this study will inspire researchers to develop new methods for field boundary segmentation and measure improvement using the FTW benchmark. We suggest benchmarking performance on the per-country test sets and reporting individual country results, the mean across all countries, or the minimum across countries (worst-case performance). Supplement Table 13 reports these metrics for the best model evaluated in this paper.
Ethics statement
Researchers, practitioners, and other users of Fields of The World must be aware of important ethical considerations raised by the digitization of field boundaries and other information from publicly accessible satellite data. Digitized field boundary data could inadvertently expose the practices and characteristics of individual land parcels, which could infringe on the privacy of local landowners who may be unaware of this digitization or its implications. There are also risks that private or public entities may use digitized field boundary data in a way that marginalizes vulnerable individuals such as smallholder farmers. Rolf et al. (2024) summarized distinct ethical concerns of machine learning applied to satellite data. In line with the recommendations of Rolf et al. (2024), we suggest that users of FTW work with local organizations and communities to build and release responsible field boundary datasets and ensure their project goals and practices align with local needs and regulations.
@@ Appendix A Supplementary Information
Figure 9 shows the distribution of (log) field area across FTW and two previous benchmark datasets.
@@
Figure 9: The distribution of (log) field area in FTW and two previous benchmark datasets.
@@
Figures 15 and 16 illustrate these predictions for a sample of 8 countries, representing both Presence/Absence and Presence-only regions (respectively). The heatmaps use three classes for prediction: Red for the Background class, Green for the Field Extent (Interior) class, and Blue for the Boundary class.
@@
The results in Tables 3, 4, and 5 of the main paper were run with one arbitrarily-chosen random seed. In supplement Tables 14-15, we report the average results across three random seeds for all test countries for the mask type experiment (Table 3 in the main paper). The standard deviation across random seeds for each experiment and test country are very small (0 or close to 0 for most metrics).
Benchmarking example
We suggest benchmarking performance on the per-country test sets and reporting individual country results, the mean across all countries, or the minimum across countries (worst-case performance). Supplement Table 13 reports these metrics for the best model evaluated in this paper (U-net with EfficientNet-b3 backbone).
Table 13: Performance metrics for U-net with EfficientNet-b3 backbone with 3-class masks, ignoring background pixels for presence-only samples. We only report recall metrics for presence-only countries.
Test country Pixel IoU Pixel precision Pixel recall Object precision Object recall
Austria 0.70 0.90 0.76 0.44 0.39
Belgium 0.75 0.92 0.80 0.57 0.58
Brazil - - 0.96 - 0.58
Cambodia 0.43 0.95 0.44 0.26 0.20
Corsica 0.48 0.79 0.55 0.21 0.17
Croatia 0.68 0.89 0.74 0.25 0.34
Denmark 0.83 0.93 0.88 0.45 0.60
Estonia 0.79 0.91 0.86 0.47 0.43
Finland 0.83 0.96 0.87 0.55 0.57
France 0.79 0.89 0.88 0.55 0.58
Germany 0.79 0.87 0.90 0.43 0.42
India - - 0.22 - 0.06
Kenya - - 0.49 - 0.10
Latvia 0.81 0.94 0.86 0.44 0.45
Lithuania 0.74 0.88 0.83 0.37 0.41
Luxembourg 0.79 0.96 0.82 0.47 0.51
Netherlands 0.75 0.92 0.80 0.53 0.45
Portugal 0.12 0.67 0.12 0.07 0.03
Rwanda - - 0.57 - 0.30
Slovakia 0.92 0.98 0.95 0.50 0.55
Slovenia 0.59 0.90 0.63 0.33 0.20
South Africa 0.79 0.89 0.88 0.51 0.55
Spain 0.74 0.96 0.76 0.36 0.20
Sweden 0.81 0.94 0.85 0.40 0.51
Vietnam 0.47 0.89 0.49 0.18 0.13
Mean 0.70 0.90 0.72 0.40 0.37
Minimum 0.12 0.67 0.12 0.07 0.03